# Sharp Eigenfunction Bounds on the Torus for large $p$

### 11 Mar 2026 v1

@@ 2.2. Proof of $\log$-loss result
These functions have the following crucial property which is trivially verified.
@@
Recall we can split our kernel as
@@
Here, $\epsilon_{0}>0$ is arbitrary. The bound for large heights can be improved to just a $\log N$ factor if $\alpha\lesssim N^{\frac{d-2}{2}-\epsilon_{0}}$. Note that the bound eq. 16 is vacuous if $d\leq 4$. The bound $|E_{\alpha}|\lesssim N^{\epsilon}\alpha^{-\frac{2(d+1)}{d-1}}$ actually holds for all level sets as a consequence of the conjecture being proven by decoupling when $p\leq 2(d+1)/(d-1)$. We can thus view this result as improving the level set bound for larger heights.
@@ 3. A sharp argument
In this section we prove Theorem 1.2.
@@ 3.1. Preliminary estimates
In this section we use different definitions for the parameters as compared to Section 2. Let $2\leq Q<N/100$ be dyadic. We also define the dyadic $2^{s}$ by $Q<2^{s}<N$. Let $\eta$ be a smooth bump function identically $1$ on $\frac{1}{4}\leq|t|\leq\frac{1}{2}$ and supported in $\frac{1}{8}\leq|t|\leq 1$. W define
@@
Note we place no restriction on $q$ other then it must be $\sim Q$. Clearly $R_{Q}\sim Q^{2}$. Our definition of $\eta$ implies
$\text{supp}\,\eta_{Q,s}=\{|t-\frac{a}{q}|\sim\frac{1}{N2^{s}},\frac{a}{q}\in R_{Q}\}.$
@@
$K_{0}\coloneq\int_{[0,1]}\prod_{j=1}^{d}G(t,x_{j})e(-\lambda t)\eta_{0}(Nt)dt,$
@@
The first piece of the kernel is supported only around $t=0$. We could alter the lower bound on $Q$ to include this portion of the domain in the definition of the $K_{Q,s}$, but we will present a simpler and distinct argument for $K_{0}$. We recall that we identify $[0,1]$ with $S^{1}$ to make sense of this definition. The cutoff function $\eta_{Q,s}$ is supported in time near rationals with denominator $\sim Q$ at a distance $\sim(N2^{s})^{-1}$. The error $K_{err}$ defined by a cutoff supported near times that cannot be well approximated by rationals of denominator $\lesssim N$.
@@
Let $d\geq 5.$ Let $2\leq p\leq\infty$. Let $K_{0}$ be as before. Then for all $F\in L^{p^{\prime}}$, the Hölder dual of $L^{p}$, we have
@@
$K_{0}(x)=\int_{[0,1]}\eta_{0}(Nt)\prod_{j=1}^{d}\sum_{k_{j}}e(x_{j}k_{j}+k_{j}^{2}t)e(-\lambda t)dt.$
We may take the Fourier transform with respect to time, as after some trivial identifications we can replace $[0,1]$ with integration over $\mathbb{R}$. Then
$K_{0}(x)=\sum_{\mathbf{k}\in\mathbb{Z}^{d}}N^{-1}\widehat{\eta_{0}}(N^{-1}(|\mathbf{k}|^{2}-\lambda))e(\mathbf{k}\cdot x).$
Recall that $\eta_{0}$ is Schwartz, and so its Fourier transform enjoys rapid decay outside of $B(0,c)$ for some $c$. By the lattice point counts, we bound the number of $\mathbf{k}$ such that $|\mathbf{k}|^{2}-\lambda=X$ with $X\lesssim N$ by $N\cdot N^{d-2}$, where there is no additional $\epsilon$ factor as $d\geq 5$. By the triangle inequality, we therefore have
@@
$\mathcal{F}(K_{0})(\mathbf{k})=\int_{\mathbb{R}^{d}}\Bigl(\int_{[0,1]}\eta_{0}(Nt)\prod_{j=1}^{d}\sum_{k}\gamma(\frac{k_{j}}{N})e(x_{j}k_{j}+k_{j}^{2}t-\lambda t)dt\Bigr)e(-\mathbf{k}\cdot x)dx,$
@@
$\mathcal{F}(K_{0})(\mathbf{k})=N^{-1}\widehat{\eta_{0}}(N^{-1}(|\mathbf{k^{2}}|-\lambda))\prod_{j=1}^{d}\gamma(\frac{k_{i}}{N}).$
@@
(28) ${\mathcal{F}(K_{0})}||_{\infty}\lesssim N^{-1}.$
@@
Our proof in dimension $2,3,4$ will also imply the sharp bound for $K_{0}*F$, but with an additional $\epsilon$ factor. Recall such a factor is necessary in $L^{p}$ estimates because of the lower bound for the number of lattice points on spheres in low dimensions has such a factor. The advantage of isolating the local contribution is that we are now comfortable proving stronger bounds for the other pieces of the kernel without fear of a contradiction, although this does not play into our proof.
@@
In contrast to the previous section, $q\geq N$. The trivial inequality and stationary phase imply as before that
@@
We need some care before we deploy this bound. Recalling (30), we will first break the domain up based on the disjoint components of $\eta_{Q,s}$, which are centered around fractions $a/q$. On each of these components $S(a,m,q)$ is constant as it does not depend on $\varphi$, and so we can sum in $a$ and $q$. Additionally, $J(x,\varphi,m,q)$ does not depend on $a$ and so has the same values on any component of the support of $\eta_{Q,s}$. This gives after a change of variables
@@
Here, $E$ is a set of measure $\sim(N2^{s})^{-1}$. As we have absorbed the sum in $a$, we now use the triangle inequality. Our bound for $\mathcal{S}(m,q)$ is independent of $m$, so much like in the proof of Lemma 2.3 we focus on the $J$ term to sum.
@@
This is a direct consequence of (2.15) in [2], taking $Q\sim 2^{s}\sim N$. Note that Proposition 3.5 would imply ${K_{err}}||_{\infty}\lesssim N^{\frac{d}{2}}$, and so the reader may wonder why are are not trying to improve (29) using cancellation like is done in cited works. The reason is subtle and will be fully elaborated on in Section 4.2 in the context of motivating computations. For now, we say we need estimates that hold for all $Q$, and the improved estimates in [4] only hold for $Q\geq N^{c}$ for small $c>0$ if $N^{\epsilon}$ errors are intolerable. As the error term is essentially the case $Q\sim N$, there is no harm having an $\epsilon$ loss there. Our pursuit of sharp estimates often limits us to only using the trivial bounds for exponential sums; finding a way around this would improve the results presented here. A similar restriction appears in the proof of the next proposition.
@@
(34) $\mathcal{F}(K_{Q,s})(\mathbf{k})\lesssim\frac{Q^{2}}{N2^{s}}.$
@@
$\mathcal{F}(K_{Q,s})(\mathbf{k})\lesssim(N2^{s})^{-1}Q^{2}.$
@@ 3.2. Proof of sharp result
We have for general $F$ that
@@
We wish to interpolate these these two estimates and sum over $Q$ and $s$. The estimate will always be largest when $2^{s}\sim N$, and we will be able to proceed when the power of $Q$ is negative, which means the estimate is dominated by the $Q\sim 1$ terms. Importantly, we have the correct powers of $N$ and $2^{s}$ without any $\epsilon$-loss which, in the end, gives us the desired sharp estimate that agrees with the local term $K_{0}$, albeit in a restricted range of $p$. Interpolation yields
@@
These functions have the following crucial property which is trivially verified.
@@
Recall we can split our kernel as
@@
The second term is handled by the estimate $|E_{\alpha}|\lesssim N^{\epsilon}\alpha^{-\frac{2(n+1)}{n-1}}$, which is a corollary of the $\ell^{2}$-decoupling theorem. For the first term we use (36). Let $p>p_{1}>2d/(d-4)$. Then
@@ 4. Exponential Sum Estimates
Recall that the generalized quadratic Gauss sums $S(a,m,q)$ are given for integers $a,m,q$ such that $(a,q)=1$ by
@@
We use a different sign convention for $m$ in this section but this has no bearing on our results. We also have the following expression which plays the role of the singular series from the circle method.
@@ 4.1. Gauss Sum Estimates
We may further assume that $(a,q)=1$, however the multiplicative nature of quadratic Gauss sums means this reduction is redundant for the bound we are trying to state. We will show that these sums exhibit square root cancellation. The $S(a,m,q)$ generalize the standard quadratic Gauss sums
@@
It is known this sum exhibits square root cancellation, which we prove here for pure primes.
@@
The assumption that $p$ is odd means that multiplication by $4$ as we sum over the cyclic groups is an isomorphism, and so we can replace it with $1$. If $\alpha\neq 0$, then by the geometric series formula the sum in $\beta$ is 0. Otherwise, it is $p$. This gives
@@
To prove an identical bound for any odd natural number $q$, one may use the imported Lemma 4.5 below. For a more elementary proof, we refer to [14]. For even numbers, there may be an additional factor of $\sqrt{2}$ if $q$ is a multiple of $4$, see Corollary 1.2.3 in [1] for a full statement.
Now we state a series of technical results which will use, along with the more general version of Lemma 4.2, to prove Proposition 4.1.
@@
Because of multiplicativity it is sufficient to show the result with $q=2^{k}$ with $k\geq 2$. We have
@@
Therefore, we can have, at most, one odd solution and one even solution. As $k$ varies in $\{0,..,2^{j}-1\}$, the value $ak^{2}+mk$ runs through each even number in $\mathbb{Z}/2^{j}\mathbb{Z}$ exactly twice by basic counting. Therefore we can write the sum as
@@
These results are enough to prove the required bound for all natural numbers $q$.
@@
for all primes $p$ nd integers $a,j$ such that $j\geq 1$ and $(a,q)=1$.
Suppose $p$ is odd. Then $p^{j}$ is as well. In what follows, denote $x^{*}$ the multiplicative inverse of $x$ modulo $p^{j}$ and $(\frac{a}{p^{j}})$ the Jacobi symbol. Direct computation gives
@@
$=\frac{1}{p^{j}}e(-4^{*}a^{*}m^{2}/p^{j})(\frac{a}{p^{j}})S(1,0,p^{j}).$
So the result follows in the case from the bound on the standard quadratic Gauss sum. Now suppose $p^{j}=2^{j}$ for some j. The result is trivial if $j=1$. If $j\geq 2$, then $2^{j}$ is divisible by $4$. If $m$ is odd, then Lemma 4.4 immediately gives the result as the sum is 0. If $j$ is even, then we can apply Lemma 4.5 with $q=2^{j}$ to get
@@
These bounds suffice for our results in Section 2. In Section 3, we sum over $a$ in the expression for $K_{Q,s}$, yielding $\mathcal{S}(m,q)$. Ostensibly the sum over $a$ gives additional cancellation, however we will argue in the following subsection this is insufficient for proving sharp results.
@@ 4.2. Difficulty in improving the range of $p$
(43) $\mathcal{S}(m,q)=S(1,0,q)^{d}\sum_{(a,q)=1}\Bigl(\frac{a}{q}\Bigr)^{d}e((-\nu/q)a^{*}-(\lambda/q)a),$
@@
The proof is standard, but we do not present it as we will not prove results in this subsection. Recall that as the Jacobi symbol takes on the values 0,1,10,1,-1, the expression for $\mathcal{S}(m,q)$ reduces to an expression involving Kloosterman sums when $d$ is even and Sailé sums when $d$ is odd. We now recall the definitions of these sums.
@@
(45) $\textbf{Sa}(\alpha,\beta,q)=\sum_{(a,q)=1}(\frac{a}{q})e(\frac{a\alpha+a^{*}\beta}{q}).$
@@
Here $\tau(q)$ is the number of divisors of $q$. As we assume $(a,q)=1$, the bound simplifies in our case to
$\textbf{Kl}(\alpha,\beta,q),\textbf{Sa}(\alpha,\beta,q)\lesssim q^{\epsilon}\sqrt{\gcd(\lambda,q)}\sqrt{q}.$
This is the bound given by (2.12) in [2]. The power of $q^{\epsilon}$ is inconsequential for our purposes. The factor $\sqrt{\gcd(\lambda,q)}$ is not. If $q|\lambda$, this means we get no cancellation in the sum and we default to using the trivial bound $\lesssim q$ from the triangle inequality.
An implicit step in [2] and [4] is the following. Restrict to $Q\leq N^{\epsilon_{0}}$ and use the triangle inequality. For $Q\geq N^{\epsilon_{0}}$, apply the above bound. The reason for this is that the additional factors of $\sqrt{(\lambda,q)}$ build up in a fatal way. If $Q$ is moderately sized, there cannot be too many $q$ that divide $\lambda$ as $\lambda\leq N^{2}$. However, for very small $q$, it may be the case that $\lambda$ has many divisors in the denominators of $R_{Q}$.
@@
This integer satisfies $X\lesssim(\log\log N)^{\log\log N}\lesssim N$. So if $X|\lambda$, we may only use the trivial estimate for every $q\sim Q$. With our interpolation, we desire a negative power of $Q$, so our estimate is dominated by the $Q\sim 1$ terms. Therefore, we must use the trivial estimates near $Q\sim 1$, which restricts the range of $p$ we can apply our argument in.
A similar phenomenon explains why we cannot improve the bound Proposition 3.7 to $Q^{1+\epsilon}$ like (14) in [4]. For small $Q$, the best we can hope for is the triangle inequality. The improved bound for moderate $Q\gtrsim N^{\epsilon_{0}}$ does not change our final bound as we are limited by the terms $Q\sim 1$.
@@
If we assume $\lambda$ does not have many divisors, we do get an improved bound. In the ideal situation, we have $\lambda$ being prime, and the techniques of this paper give the sharp results for $p>\frac{2(d+1)}{d-3}$ for $d\geq 5$ by using the full square root cancellation in Kloosterman and Sailé sums. Intermediate bounds can be proven based on the structure of the divisors of $\lambda$, but as we are interested in the strongest bound for the full set of eigenvalues we do not record them here.

### 06 Apr 2026 v2

@@ 2.2. Proof of $\log$-loss result
These functions have the following crucial property, which is trivially verified.
@@
Recall, we can split our kernel as
@@
Here, $\epsilon_{0}>0$ is arbitrary. The bound for large heights can be improved to just a $\log N$ factor if $\alpha\lesssim N^{\frac{d-2}{2}-\epsilon_{0}}$. Note that the bound (16) is vacuous if $d\leq 4$. The bound $|E_{\alpha}|\lesssim N^{\epsilon}\alpha^{-\frac{2(d+1)}{d-1}}$ actually holds for all level sets as a consequence of the conjecture being proven by decoupling when $p\leq 2(d+1)/(d-1)$. We can thus view this result as improving the level set bound for larger heights.
@@ 3. A sharp argument
In this section, we prove Theorem 1.2.
@@ 3.1. Preliminary estimates
In this section we use different definitions for parameters as compared to Section 2. Let $2\leq Q<N/100$ be dyadic. We also define the dyadic $2^{s}$ by $Q<2^{s}<N$. Let $\eta$ be a smooth bump function identically $1$ on $\frac{1}{4}\leq|t|\leq\frac{1}{2}$ and supported in $\frac{1}{8}\leq|t|\leq 1$. We define
@@
Note we place no restriction on $q$ other then it must be $\sim Q$. Clearly, $R_{Q}\sim Q^{2}$. Our definition of $\eta$ implies
$\text{supp}\,\eta_{Q,s}=\{t:|t-\frac{a}{q}|\sim\frac{1}{N2^{s}},\frac{a}{q}\in R_{Q}\}.$
@@
$K_{0}\coloneq\int_{[0,1]}\prod_{j=1}^{d}G(t,x_{j})e(-\lambda t)\eta_{0}(N^{2}t)dt,$
@@
The first piece of the kernel is supported only around $t=0$. We could alter the lower bound on $Q$ to include this portion of the domain in the definition of the $K_{Q,s}$, but we will present a simpler and distinct argument for $K_{0}$. We recall that we identify $[0,1]$ with $S^{1}$ to make sense of this definition and the dyadic arcs around 0. The cutoff function $\eta_{Q,s}$ is supported in time near rationals with denominator $\sim Q$ at a distance $\sim(N2^{s})^{-1}$. The error $K_{err}$ is defined by a cutoff supported near times that cannot be well approximated by rationals of denominator $\lesssim N$.
@@
Let $d\geq 5$ and $2\leq p\leq\infty$. Let $K_{0}$ be as before. Then, for all $F\in L^{p^{\prime}}$, the Hölder dual of $L^{p}$, we have
@@
$K_{0}(x)=\int_{[0,1]}\eta_{0}(N^{2}t)\prod_{j=1}^{d}\sum_{k_{j}}\gamma(k_{j}/N)e(x_{j}k_{j}+k_{j}^{2}t)e(-\lambda t)dt.$
We may take the Fourier transform with respect to time as, after some trivial identifications, we can replace $[0,1]$ with integration over $\mathbb{R}$. Then
$K_{0}(x)=\sum_{\mathbf{k}\in\mathbb{Z}^{d}}N^{-2}\widehat{\eta_{0}}(N^{-2}(|\mathbf{k}|^{2}-\lambda))e(\mathbf{k}\cdot x).$
Recall that $\eta_{0}$ is Schwartz, and so its Fourier transform enjoys rapid decay outside of $B(0,c)$ for some $c$. By the lattice point counts, we bound the number of $\mathbf{k}$ such that $|\mathbf{k}|^{2}-\lambda=X$ with $X\lesssim N^{2}$ by $N^{2}\cdot N^{d-2}$, where there are no additional $\epsilon$ factors as $d\geq 5$. By the triangle inequality, we therefore have
@@
$\mathcal{F}(K_{0})(\mathbf{k})=\int_{\mathbb{R}^{d}}\Bigl(\int_{[0,1]}\eta_{0}(N^{2}t)\prod_{j=1}^{d}\sum_{k}\gamma(\frac{k_{j}}{N})e(x_{j}k_{j}+k_{j}^{2}t-\lambda t)dt\Bigr)e(-\mathbf{k}\cdot x)dx,$
@@
$\mathcal{F}(K_{0})(\mathbf{k})=N^{-2}\widehat{\eta_{0}}(N^{-2}(|\mathbf{k^{2}}|-\lambda))\prod_{j=1}^{d}\gamma(\frac{k_{i}}{N}).$
@@
(28) ${\mathcal{F}(K_{0})}||_{\infty}\lesssim N^{-2}.$
@@
Our proof in dimension $2,3,4$ will also imply the sharp bound for $K_{0}*F$, but with an additional $\epsilon$ factor. Recall such a factor is necessary in $L^{p}$ estimates because the lower bound for the number of lattice points on spheres in low dimensions has such a factor. The advantage of isolating the local contribution is that we are now comfortable proving stronger bounds for the other pieces of the kernel without fear of a contradiction, although this does not play into our proof.
@@
In contrast to the previous section, $q\leq N$ here. The trivial inequality and stationary phase imply as before that
@@
We need some care before we deploy this bound. Recalling (30), we will first break the domain up based on the disjoint components of $\eta_{Q,s}$, which are centered around fractions $a/q$. On each of these components, $S(a,m,q)$ is constant as it does not depend on $\varphi$, and so we can sum in $a$ and $q$. Additionally, $J(x,\varphi,m,q)$ does not depend on $a$ and so has the same values on any component of the support of $\eta_{Q,s}$. This gives after a change of variables
@@
Here, $E$ is a set of measure $\sim(N2^{s})^{-1}$. As we have absorbed the sum in $a$, we now use the triangle inequality. Our bound for $\mathcal{S}(m,q)$ is independent of $m$, so, much like in the proof of Lemma 2.3, we focus on the $J$ term to sum.
@@
This is a direct consequence of (2.15) in [2], taking $Q\sim 2^{s}\sim N$. Note that Proposition 3.5 would imply ${K_{err}}||_{\infty}\lesssim N^{\frac{d}{2}}$. Such improvements are not obvious for (29) using cancellation like is done in cited works because of our intolerance of $\epsilon$-losses. This will be discussed more a the end of Section 4, but the improved estimates in [4] only hold for $Q\geq N^{c}$ for small $c>0$. As the error term is essentially the case $Q\sim N$, there is no harm having an $\epsilon$-loss in $N$ or $2^{s}$ there. The main term in $Q\sim 1$. Our pursuit of sharp estimates often limits us to only using the trivial bounds for exponential sums; finding a way around this would improve the results presented here. A similar restriction appears in the proof of the next proposition.
@@
(34) $|\mathcal{F}(K_{Q,s})(\mathbf{k})|\lesssim\frac{Q^{2}}{N2^{s}}.$
@@
$|\mathcal{F}(K_{Q,s})(\mathbf{k})|\lesssim(N2^{s})^{-1}Q^{2}.$
@@ 3.2. Proof of sharp result
Using the estimates proven in the previous subsection and basic inequalities, we have, for general $F$, that
@@
We wish to interpolate these these two estimates and sum over $Q$ and $s$. The estimate will always be largest when $2^{s}\sim N$, and we will be able to proceed when the power of $Q$ is negative, which means the estimate is dominated by the $Q\sim 1$ terms. Importantly, we have the correct powers of $N$ and $2^{s}$ without any $\epsilon$-loss which, in the end, give us the desired sharp estimate that agrees with the local term $K_{0}$, albeit in a restricted range of $p$. Interpolation yields
@@
These functions have the following crucial property, which is trivially verified.
@@
Recall, we can split our kernel as
@@
The second term is handled by the estimate $|E_{\alpha}|\lesssim N^{\epsilon}\alpha^{-\frac{2(n+1)}{n-1}}$, which is a corollary of the $\ell^{2}$-decoupling theorem. For the first term, we use (36). Let $p>p_{1}>2d/(d-4)$. Then
@@ 4. Exponential Sum Estimates
Recall that the generalized quadratic Gauss sums $S(a,m,q)$ are given, for integers $a,m,q$ such that $(a,q)=1$, by
@@
We use a different sign convention for $m$ in this section, but this has no bearing on our results. We also have the following expression, which plays the role of the singular series from the circle method.
@@ 4.1. Gauss Sum Estimates
We may further assume that $(a,q)=1$, however the multiplicative nature of quadratic Gauss sums means this reduction is redundant for the bound we are trying to state. We will show that these sums exhibit square root cancellation. The $S(a,m,q)$ generalize the standard quadratic Gauss sums.
@@
It is known this sum exhibits square root cancellation, which we prove here for odd primes.
@@
The assumption that $p$ is odd means that multiplication by $4$ as we sum over the cyclic groups is an isomorphism, and so we can replace it with $1$. If $\alpha\neq 0$, then, by the geometric series formula, the sum in $\beta$ is 0. Otherwise, it is $p$. This gives
@@
To prove an identical bound for any odd natural number $q$, one may use the imported Lemma 4.5 below. For a more elementary proof, we refer to [14]. The case $p=2$ is trivial. For general even numbers, there may be an additional factor of $\sqrt{2}$ if $q$ is a multiple of $4$. See Corollary 1.2.3 in [1] for a full statement. With this imported estimate, we have Lemma 4.2 for all integers after proving the next lemma.
We now also state a series of technical results which we will use, along with the more general version of Lemma 4.2, to prove Proposition 4.1.
@@
Because of multiplicativity, it is sufficient to show the result with $q=2^{k}$ and $k\geq 2$. We have
@@
Therefore, we can have, at most, one odd solution and one even solution mod $2^{j}$. As $k$ varies in $\{0,...,2^{j}-1\}$, the value $ak^{2}+mk$ runs through each even number in $\mathbb{Z}/2^{j}\mathbb{Z}$ exactly twice by basic counting. We can write the sum as
@@
These results are enough to prove Proposition 4.1 for all natural numbers $q$.
@@
for all primes $p$ and integers $a,j$ such that $j\geq 1$ and $(a,q)=1$. In the case $p=2$, we allow a multiplied constant.
Suppose $p$ is odd. Then $p^{j}$ is as well. In what follows, denote $x^{*}$ the multiplicative inverse of $x$ modulo $p^{j}$ and $(\frac{a}{p^{j}})$ the Jacobi symbol. Direct computation and completing the square gives
@@
$=e(-4^{*}a^{*}m^{2}/p^{j})(\frac{a}{p^{j}})S(1,0,p^{j}).$
So the result follows in the case from the bound on the standard quadratic Gauss sum. Now suppose $p^{j}=2^{j}$ for some j. Recall the bound for $S(1,0,2^{j})$ follows from Corollary 1.2.3 in [1].
The result is trivial if $j=1$. If $j\geq 2$, then $2^{j}$ is divisible by $4$. If $m$ is odd, then Lemma 4.4 immediately gives the result as the sum is 0. If $m$ is even, then we can apply Lemma 4.5 with $q=2^{j}$ to get
@@
These bounds suffice for our results in Section 2. In Section 3, we sum over $a$ in the expression for $K_{Q,s}$, yielding $\mathcal{S}(m,q)$. Ostensibly, the sum over $a$ gives additional cancellation, however, it is difficult to leverage this to prove sharp results.
@@ 4.2. Difficulty in improving the range of $p$
(43) $\mathcal{S}(m,q)=S(1,0,q)^{d}\sum_{(a,q)=1}\Bigl(\frac{a}{q}\Bigr)^{d}e((\nu/q)a^{*}-(\lambda/q)a),$
@@
The proof is standard, but we do not present it as we will not prove results in this subsection. As the Jacobi symbol takes on the values 0,1,10,1,-1, the expression for $\mathcal{S}(m,q)$ reduces to an expression involving Kloosterman sums when $d$ is even and Sailé sums when $d$ is odd. We now recall the definitions of these sums.
@@
(45) $\textbf{Sa}(\alpha,\beta,q)\coloneq\sum_{(a,q)=1}(\frac{a}{q})e(\frac{a\alpha+a^{*}\beta}{q}).$
@@
Here $\tau(q)$ is the number of divisors of $q$. In our case we have $\alpha=-\lambda$ and $\beta=\nu$. As we assume $(a,q)=1$, the bound simplifies in our case to
$\textbf{Kl}(-\lambda,\nu,q),\textbf{Sa}(-\lambda,\nu,q)\lesssim q^{\epsilon}\sqrt{\gcd(\nu,\lambda,q)}\sqrt{q}.$
This is the bound stated at the end of Section 3 in [4]. The power of $q^{\epsilon}$ is inconsequential for our purposes. The factor $\sqrt{\gcd(\lambda,q)}$ is not. If $q|\lambda$, this means we get no cancellation in the sum using this bound, and we default to using the trivial bound $\lesssim q$ from the triangle inequality. However, both the Kloosterman and Sailé sums simplify in this case to sums with at least square root cancellation, although a different argument must be presented. There are additional cases to consider, however. $\lambda$ and $q$ could have a large common factor but $q$ not divide $\lambda$. In this case, we simplify to a sum that is not classical and the analysis is more complicated. There is also an additional summation over $m$ that may assist in proving better bounds. We intend to explore this in future work to prove both sharp bounds and bounds with an $\epsilon$-loss.
A difficulty with improving the bound in Proposition 3.7 is the following. An implicit step in [2] and [4] is to restrict to $Q\lesssim N^{\epsilon_{0}}$ and use the triangle inequality. One can prove an improved bound, replacing $Q^{2}$ by $Q$ in our proposition, for $Q\gtrsim N^{\epsilon_{0}}$ using essentially the same proof as for Proposition 2.5 and an additional step using a mean zero normalization. For small $Q$, the best we can hope for is the triangle inequality. The improved bound for moderate $Q\gtrsim N^{\epsilon_{0}}$ does not change our final bound as we are limited by the terms $Q\sim 1$.
@@
This integer satisfies $X\lesssim(\log\log N)^{\log\log N}\lesssim N$. The same approach we used to estimate (18) will not work as we could have many factors that divide $l$. Therefore, we must use the trivial estimates near $Q\sim 1$, which restricts the range of $p$ we can apply our argument in.
@@
If we assume $\lambda$ does not have many divisors, we do get an improved bound. In the ideal situation, we have $\lambda$ being prime, and the techniques of this paper give the sharp results for $p>\frac{2(d+1)}{d-3}$ for $d\geq 5$, by using the full square root cancellation in Kloosterman and Sailé sums. Intermediate bounds can be proven based on the structure of the divisors of $\lambda$, but as we are interested in the strongest bound for the full set of eigenvalues we do not record them here.